the square of a number decreased by the square of one half the number is 108. find the number.
Answer by checkley71(8403)

x~2-(x/2)~2=108 or x2-x~2/4=108 or (4x~2-x~2)/4=108 or 3x~2/4=108 or
3x~2=108 or x~2=144 or x=sqrt144 or x=12
proof 12~2-(12/2)~2=108 OR 144-(144/4)=108 OR 144-36=108 OR 108=108